# Identification of Single-Treatment Effects in Factorial Experiments

### v2

@@ 4 Simulation
To illustrate the proposed method, I simulate two examples described in the Supplementary Material. The first example involves a structural model with no explicit constraints, while the second one is generated using probabilities over principal strata to enforce both monotonicity and no interaction. Then I generate firstly for each model an example dataset with size equal to $1,000$, both for the observational and for the factorial case (total of $2,000$).
Previously, I discussed identification results tailored to scenarios where researchers have access to population quantities. Here, I delve into the challenge of estimating bounds when dealing with finite samples. In such cases, an additional hurdle arises as it’s not straightforward to estimate confidence intervals for quantities derived from linear programs. Chickering and Pearl, (1996) utilize Bayesian inference via Gibbs’ sampling. Duarte et al., (2023) consider two distinct methods for calculating bounds using autobounds: one based on Kullback-Leibler divergence of Bernoulli distribution, and another utilizing multivariate Gaussian limiting distribution of a Dirichlet. Levis et al., (2023) compute influence functions when incorporating covariates. I remain neutral regarding the proposed solutions. Since the solution provided pertains to identification rather than inference, any of these methods can be employed if the underlying assumptions are justified. To illustrate the bounds with the generated data, I employ bootstrap. So, I generate $1000$ resamples of the generated initial sample and calculate 95% confidence bounds both for the lower bounds and for the upper bounds.
@@
Example 1 was simulated nonparametrically using structural equations, without assuming non-interactivity and monotonicity. The model produced an ATE of 0.24. Two sets of results were estimated: firstly, nonparametric bounds assuming only factorial data, and secondly, nonparametric bounds using both factorial and observational data. These results are depicted in the left panel of Figure 6. The estimated bounds (blue lines) encompass the true ATE value (black lines). The version incorporating both distributions yields bounds superior to the second version. Additionally, I estimated a 95% confidence region (in red) using resampling techniques. In this particular example, one cannot distinguish that the real ATE is different from 0.
@@
Figure 6: Estimated bounds using simulated data from examples 1 and 2. The blue bars represent the estimated bounds, while the red lines indicate the 95% confidence region for the bounds calculated using bootstrap. The dotted black line shows the true value of the Average Treatment Effect (ATE) for each example: 0.24 for the first and 0.58 for the second. Example 2 explicitly contains only no interactive and monotonic units. Therefore, the analysis for the first model only shows nonparametric bounds. As expected, both bounds cover the real ATE, but those incorporating both factorial and observational data are sharper than those with factorial data alone. For the second example, we estimated three versions: one assuming monotonicity, another assuming a maximum proportion of 0.15 non-interactive units, and a nonparametric version. As expected, the bounds in the nonparametric version are lower than in the other two. In other words, justified restrictive assumptions lead to a quantitative gain in causal information.
@@
In example 2, I explicitly made assumptions of monotonicity and no interaction, with a true ATE of 0.58. This justifies estimating bounds based on these assumptions. As demonstrated, assuming no interaction can lead to point identification results. Rather than adopting this approach, I considered a scenario where one might wish to examine the sensitivity of the no-interaction assumption. In Figure 5, I illustrate bounds conditioned on the maximum proportion of non-interactive units permitted in the model. For instance, if no constraint is imposed and full interactions are allowed, then nonparametric bounds are obtained. All other cases lie between the full interaction and non-interactive point identifiable scenarios. After analyzing the graph, suppose one believes that a maximum of 15% of non-interactive units is reasonable; in that case, the bounds can be evaluated accordingly, along with confidence intervals. This result is depicted in the right panel of Figure 6. Additionally, bounds for the case assuming only monotonicity were included, showing marked improvements compared to the nonparametric scenario. However, all these cases, even when utilizing only the factorial distribution, provide an opportunity to estimate bounds greater than 0 (which can be signed) and encompass the real ATE.
@@ 5 Discussion
The paper’s findings highlight that while factorial experiments can significantly reduce the costs associated with estimating multiple treatment effects, they also introduce complexities when estimating single-treatment effects. Specifically, when considering interactions, these effects become non-point-identifiable, and the best approach is to estimate sharp bounds. To address this, I provided bounds for both the nonparametric case and the case assuming monotonicity, aiming to assist researchers in estimating effects under these conditions. Moreover, I outlined a procedure for sensitivity analysis, enabling researchers to justify other assumptions, such as non-interaction, or to test the sensitivity of results to a maximum proportion of non-interactive units. For instance, by limiting interactive units to at most 15%, researchers can still derive meaningful bounds.
@@ A.1 Example 1
From this example, one can calculate quantities such as the real ATE, which is equal to $0.24$. However, from the factorial experiment, obtained quantities refer not to interventions exclusively on $A$, but on $A$ and $B$ at the same time. In particular, the difference of the expected value of $Y$ given interventions on $A=a_{1}$ and $A=a_{0}$, given that we also had an intervention on $B=b$ are respectively, $-0.55$ and $0.042$, for the cases where $B=b_{0}$ and $B=b_{1}$. Surprisingly, one finds out that using the estimator described by the first researcher – which averages those differences according to a uniform distribution, i.e. $P(B=b_{1})=P(B=b_{0})=0.5$ – then the estimated value will be equal to $-0.25$. Alternatively, using the estimator described by the second researcher – which employs the population distribution of $B$ – the estimated value will be equal to $-0.07$. Therefore, contrarily to the real positive value of the ATE, we obtained two negative estimates. In fact, one can realize that there is no distribution over $B$ that can combine the factorial differences into the real ATE. This implies any attempt to use factorial experiments to answer queries concerning marginal treatments will necessarily be biased. I refer to the bias induced when extrapolating from factorial experiments to in-the-wild effects of isolated treatments as F-bias.
@@
In particular, all non-monotonic and non-interactive strata were set to have probability equal to 00. So only probability over the strata $y_{0000},y_{0101},y_{0011},y_{1111}$ were allowed. This model yields an ATE $0.58$. Eventual calculations are provided in the attached python code.
@@ Appendix B Non-identifiability results
Consider treatments $A$ and $B$, along with an outcome $Y$. Assume that both $A$ and $B$ independently cause $Y$, without causing each other. Additionally, suppose the effects of both $A$ and $B$ on $Y$ are monotonic, meaning $\Pr(Y_{a_{0}}=y_{1},Y_{a_{1}}=y_{0})=\Pr(Y_{b_{0}}=y_{1},Y_{b_{1}}=y_{0})=0$. Under these conditions, without imposing any other assumptions, such as absence of confounding between $A$, $B$, and $Y$, the probability distribution $\Pr(Y_{a})$, representing the outcome under different levels of treatment $A$, is not identifiable solely from the joint distribution $\Pr(Y_{a,b})$ and $\Pr(Y,A,B)$. Furthermore, the Average Treatment Effect (ATE), defined as the difference in probabilities between different treatment levels, $\Pr(Y_{a}=y)-\Pr(Y_{a^{\prime}}=y)$, remains unidentifiable. Finally, these quantities are not identifiable even if we relax the assumption of monotonicity or if we only have observational or factorial data alone.
@@
Now, let the outcome $Y$ for model 00, $Y_{M_{0}}$, be determined by:
$Y_{M_{0}}=\begin{cases}\max\{A,B\},&\text{ if }U_{ay}=U_{by}=0\ \text{ and }\ U_{y}=2,\\ B,&\text{ if }U_{ay}=U_{by}=0\ \text{ and }\ U_{y}=1,\\ 1,&\text{ otherwise }\end{cases}$
@@
$Y_{M_{1}}=\begin{cases}\max\{A,B\},&\text{ if }U_{ay}=U_{by}=0\ \text{ and }\ U_{y}\neq 0,\\ \max\{1-A,B\},&\text{ if }U_{ay}=0,U_{by}=1\ \text{ and }\ U_{y}\neq 0,\\ 1,&\text{ otherwise }\end{cases}$
@@
For data generated by model $M_{0}$, if $A$ or $B$ equals $1$, then $Y$ must also equal $1$, implying probabilities of 00 for scenarios where $Y=0$ under those conditions and $0.25$ for scenarios where $Y=1$ under the same conditions. When $A=B=0$, there are three possibilities: if $U_{y}=2$ or $U_{y}=1$, then $Y=0$, and if $U_{y}=0$, then $Y=1$. This yields probabilities of $3/5$ for $Y=0$ and $2/5$ for $Y=1$.
For data generated by model $M_{1}$, if $A=1$, then $Y$ must equal $1$, resulting in probabilities of 00 for scenarios where $Y=0$ and $0.25$ for scenarios where $Y=1$. If $A=0$ and $B=1$, then $Y$ is always $1$, independently of $U_{y}$, leading to probabilities of 00 for $Y=0$ and $0.25$ for $Y=1$. When $U_{ay}=U_{by}=0$, there are two possibilities: if $U_{y}=2$ or $U_{y}=1$, then $Y=0$, and if $U_{y}=0$, then $Y=1$, resulting in probabilities of $3/5$ for $Y=0$ and $2/5$ for $Y=1$.
@@
In both models $M_{0}$ and $M_{1}$, when $B$ is forced to be $1$, $Y$ is always $1$. If $B$ is forced to be 00, and $A$ is also forced to be 00, then $Y$ is 00 when $U_{ay}=U_{by}=0$ and $U_{y}\neq 0$, resulting in probabilities of $0.15$ for $Y=0$ and $0.85$ for $Y=1$. If $A$ is forced to be $1$, then $Y$ is 00 when $U_{ay}=U_{by}=0$ and $U_{y}=1$, leading to probabilities of $0.1$ for $Y=0$ and $0.9$ for $Y=1$.
@@
Additionally, both models satisfy monotonicity with respect to $A$ and $B$. There are no cases where $Y_{a=1}=0$ and $Y_{a=0}=1$ in model $M_{1}$, and in model $M_{0}$, the only such case occurs when $Y=B$, which remains unchanged if $A$ is set to 00. Similarly, there are no cases where $Y_{b=1}=0$ and $Y_{b=0}=1$ due to the nature of $Y$ in both models.
@@
When $A$ is set to 00 regardless of $B$ in both models $M_{0}$ and $M_{1}$, only one scenario results in $Y=0$: when $U_{ay}=U_{by}=0$ and $U_{y}\neq 0$. This yields $\Pr_{M_{0}}(Y_{A=0}=0)=\Pr_{M_{1}}(Y_{A=0}=0)=0.15$. However, when $A$ is forced to be $1$, in model $M_{0}$, $Y$ consistently equals $1$ across all scenarios. Yet, in model $M_{1}$, there’s a case where $Y=0$ when $U_{ay}=U_{by}=0$ and $U_{y}=1$, with a probability of $0.1$. Thus, $\Pr_{M_{0}}(Y_{A=1}=1)=1$, different from $\Pr_{M_{1}}(Y_{A=1}=1)=0.9$. Consequently, $ATE_{M_{0}}=0.15$, in contrast to $ATE_{M_{1}}=0.05$. This disparity demonstrates that both $\Pr(Y_{A=1})$ and the ATE lack point identifiability.
Table 4: Calculated probabilities for models $M_{0}$ and $M_{1}$ as presented in the counterexample detailed in Theorem B.1. Despite both models being monotonic, they agree on all observational and factorial (2-treatment) quantities. However, they diverge regarding $\Pr(Y_{A=1})$, making not only this probability but also the Average Treatment Effect (ATE) non-identifiable
@@
$Y_{M_{0}}=\begin{cases}0,\text{ if }&U_{aby}=1,U_{y}=0\\ 1,\text{ if }&U_{aby}=1,U_{y}=1\\ B,\text{ if }&U_{aby}=0,U_{y}=0\text{ or }U_{aby}=2,U_{y}\neq 0\\ A,\text{ otherwise }&\end{cases}$
@@
$Y_{M_{1}}=\begin{cases}0,\text{ if }&U_{aby}=1,U_{y}=0\\ 1,\text{ if }&U_{aby}=1,U_{y}=1\\ B,\text{ if }&U_{aby}=2,\\ A,\text{ otherwise }&\end{cases}$
@@ Appendix C Parametric Identification of Single-Treatment Estimands Using Factorial Data
Theorem C.1 (Additive separability is sufficient for identifiability of the ATE).
@@
Suppose $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta_{2}b_{i}+\beta_{3}a_{i}b_{i}+u_{i}$ with unobserved $u_{i}$ is a model of the real-world, discrete $b_{i}$ and $a_{i}$, and assume statistical dependence between $a_{i}$ and $b_{i}$ and $u_{i}$. Then the the expected outcome of an intervention $A=a$ is identifiable using the observational quantity $\Pr(B=b)$ and the factorial distribution $\operatorname{\mathbb{E}}(Y_{a,b})$ by:
@@
Suppose $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta_{2}b_{i}+u_{i}$ with unobserved $u_{i}$ is a model of the real-world and assume statistical dependence between $a_{i}$ and $b_{i}$ and $u_{i}$. Then the ATE of $A$ on $Y$ is identifiable using factorial data by the quantity in equation 4 as that is a case of additive separability (Theorem C.1). Additionally, linearity is also a case of interactive linearity (Theorem C.2), so single-treatment quantities are identifiable by Equation 5. Finally, observational data alone is insufficient for identifiability given the standard omitted variable bias (unobserved confounding).

### v3

@@ 4 Simulation
To illustrate the proposed method, I simulate two examples described in the Supplementary Material. The first example involves a structural model with no explicit constraints, while the second one is generated using probabilities over principal strata (Greenland and Robins, 1986; Frangakis and Rubin, 2002) to enforce both monotonicity and no interaction. Then I generate firstly for each model an example dataset with size equal to $1,000$, both for the observational and for the factorial case (total of $2,000$).
Previously, I discussed identification results tailored to scenarios where researchers have access to population quantities. Here, I discuss the challenge of estimating bounds when dealing with finite samples. In such cases, an additional hurdle arises as it’s not straightforward to estimate confidence intervals for extremas of distributions. Firstly, the main issue is that the bounds are not smooth functions of the underlying distributions, which complicates the application of standard asymptotic theory. It is known that if one maximizes (or minimizes) over many candidate functions that are smooth, the resulting function may not be smooth. This eschews the use of both delta method and bootstrap. Secondly, in practice, it is often assumed that the if one knows which candidate is the correct one in the population of interest or at least approximately knows it, then one can use standard asymptotic theory. However, in finite samples, the selected candidate may differ from the population one with non-negligible probability, leading to incorrect inference. This issue has been explored extensively in the econometrics literature, notably by Andrews and Soares (2010); Bugni (2010); Canay (2010); Chernozhukov et al. (2013).
Several solutions have been proposed to address this challenge. For simplicity, I employ Bayesian Dirichlet posteriors to estimate the underlying distributions (Richardson et al., 2011; Marden et al., 2018). Specifically, I assume a Dirichlet prior with parameters equal to 1 for each cell of the joint distribution. After observing the data, I compute the posterior distribution, which is also Dirichlet, with parameters equal to the observed counts plus 1. Differently from bootstrap, which resamples the data from a multinomial distribution with parameters equal to the empirical frequencies, the Dirichlet posterior can be interpreted as sampling from a multinomial distribution with parameters equal to the smoothed frequencies. Despite not having frequentist guarantees, this approach is valid from a Bayesian perspective.
From the Dirichlet posterior distribution, I sample $1000$ times the underlying distributions and compute the bounds for each sample. So I calculate 95% confidence bounds both for the lower bounds and for the upper bounds.
@@
Example 1 was simulated nonparametrically using structural equations, without assuming non-interactivity and monotonicity. The model produced an ATE of 0.24. Two sets of results were estimated: firstly, nonparametric bounds assuming only factorial data, and secondly, nonparametric bounds using both factorial and observational data. These results are depicted in the left panel of Figure 6. The estimated bounds (blue lines) encompass the true ATE value (black lines). The version incorporating both distributions yields bounds superior to the second version. Additionally, I estimated a 95% confidence region (in red) using resampling techniques. In this particular example, one cannot rule out that the real ATE is different from 0.
@@
Figure 6: Estimated bounds using simulated data from examples 1 and 2. The blue bars represent the estimated bounds, while the red lines indicate the 95% credible region for the bounds calculated using Dirichlet posteriors. The dotted black line shows the true value of the ATE for each example: 0.24 for the first and 0.58 for the second. Example 2 explicitly contains only no interactive and monotonic units. Therefore, the analysis for the first model only shows nonparametric bounds. As expected, both bounds cover the real ATE, but those incorporating both factorial and observational data are sharper than those with factorial data alone. For the second example, we estimated three versions: one assuming monotonicity, another assuming a maximum proportion of 0.15 non-interactive units, and a nonparametric version. As expected, the bounds in the nonparametric version are lower than in the other two. In other words, justified restrictive assumptions lead to a quantitative gain in causal information.
@@
In example 2, I explicitly made assumptions of monotonicity and no interaction, with a true ATE of 0.58. This justifies estimating bounds based on these assumptions. As demonstrated, assuming no interaction can lead to point identification results. Rather than adopting this approach, I considered a scenario where one might wish to examine the sensitivity of the no-interaction assumption. In Figure 5, I illustrate bounds conditioned on the maximum proportion of non-interactive units permitted in the model. For instance, if no constraint is imposed and full interactions are allowed, then nonparametric bounds are obtained. All other cases lie between the full interaction and non-interactive point identifiable scenarios. After analyzing the graph, suppose one believes that a maximum of 15% of non-interactive units is reasonable; in that case, the bounds can be evaluated accordingly, along with credible intervals. This result is depicted in the right panel of Figure 6. Additionally, bounds for the case assuming only monotonicity were included, showing marked improvements compared to the nonparametric scenario. However, all these cases, even when utilizing only the factorial distribution, provide an opportunity to estimate bounds greater than 0 (which can be signed) and encompass the real ATE.
@@ 5 Discussion
The paper’s findings highlight that while factorial experiments can significantly reduce the costs associated with estimating multiple treatment effects, they also introduce complexities when extrapolating to single-treatment scenarios where the other treatments are not manipulated. Specifically, when considering interactions, these effects become non-recoverable, without imposing strong assumptions.
Recognizing the challenges in justifying such assumptions, and a practical approach is to estimate sharp bounds. To address this, I provided bounds for both the nonparametric case and the case assuming monotonicity, aiming to assist researchers in estimating effects under these conditions. Moreover, I outlined a procedure for sensitivity analysis, allowing researchers to explore other assumptions, such as non-interaction, or to test the sensitivity of results to a maximum proportion of non-interactive units. For instance, by limiting interactive units to at most 15%, researchers can still derive meaningful bounds. That shows how useful information on external validity can be extracted from factorial experiments.
@@ A.1 Example 1
From this example, one can calculate quantities such as the real ATE, which is equal to $0.24$. However, the factorial experiment yields quantities tied to interventions on both $A$ and $B$ at the same time. In particular, the difference of the expected value of $Y$ given interventions on $A=a_{1}$ and $A=a_{0}$, given that we also had an intervention on $B=b$ are respectively, 0.55-0.55 and $0.042$, for the cases where $B=b_{0}$ and $B=b_{1}$. Using the estimator described by the first researcher – which averages those differences according to a uniform distribution, i.e. $P(B=b_{1})=P(B=b_{0})=0.5$ – yields an estimate of 0.25-0.25. Alternatively, using the estimator described by the second researcher – which employs the population distribution of $B$ – yields an estimate of 0.07-0.07. In this example, both estimates are negative despite the positive ATE. There is no distribution over $B$ that combines the factorial differences into the real ATE in this setting. This suggests that attempts to use factorial experiments to answer queries concerning marginal treatments can be biased. I refer to the bias induced when extrapolating from factorial experiments to in-the-wild effects of isolated treatments as F-bias.
@@
In particular, all non-monotonic and non-interactive strata were set to have probability equal to 0. So only probability over the strata $y_{0000},y_{0101},y_{0011},y_{1111}$ were allowed. This model yields an ATE $0.58$. Eventual calculations are provided in the attached python code.
@@ Appendix B Non-identifiability results
Consider treatments $A$ and $B$, along with an outcome $Y$. Assume that both $A$ and $B$ independently cause $Y$, without causing each other. Additionally, suppose the effects of both $A$ and $B$ on $Y$ are monotonic, meaning $\Pr(Y_{a_{0}}=y_{1},Y_{a_{1}}=y_{0})=\Pr(Y_{b_{0}}=y_{1},Y_{b_{1}}=y_{0})=0$. Under these conditions, without imposing any other assumptions, such as absence of confounding between $A$, $B$, and $Y$, the probability distribution $\Pr(Y_{a})$, representing the outcome under different levels of treatment $A$, is not identifiable solely from the joint distribution $\Pr(Y_{a,b})$ and $\Pr(Y,A,B)$. Furthermore, the ATE, defined as the difference in probabilities between different treatment levels, $\Pr(Y_{a}=y)-\Pr(Y_{a^{\prime}}=y)$, remains unidentifiable. Finally, these quantities are not identifiable even if we relax the assumption of monotonicity or if we only have observational or factorial data alone.
@@
Now, let the outcome $Y$ for model 0, $Y_{M_{0}}$, be determined by:
[…]
@@
[…]
@@
For data generated by model $M_{0}$, if $A$ or $B$ equals $1$, then $Y$ must also equal $1$, implying probabilities of 0 for scenarios where $Y=0$ under those conditions and $0.25$ for scenarios where $Y=1$ under the same conditions. When $A=B=0$, there are three possibilities: if $U_{y}=2$ or $U_{y}=1$, then $Y=0$, and if $U_{y}=0$, then $Y=1$. This yields probabilities of $3/5$ for $Y=0$ and $2/5$ for $Y=1$.
For data generated by model $M_{1}$, if $A=1$, then $Y$ must equal $1$, resulting in probabilities of 0 for scenarios where $Y=0$ and $0.25$ for scenarios where $Y=1$. If $A=0$ and $B=1$, then $Y$ is always $1$, independently of $U_{y}$, leading to probabilities of 0 for $Y=0$ and $0.25$ for $Y=1$. When $U_{ay}=U_{by}=0$, there are two possibilities: if $U_{y}=2$ or $U_{y}=1$, then $Y=0$, and if $U_{y}=0$, then $Y=1$, resulting in probabilities of $3/5$ for $Y=0$ and $2/5$ for $Y=1$.
@@
In both models $M_{0}$ and $M_{1}$, when $B$ is forced to be $1$, $Y$ is always $1$. If $B$ is forced to be 0, and $A$ is also forced to be 0, then $Y$ is 0 when $U_{ay}=U_{by}=0$ and $U_{y}\neq 0$, resulting in probabilities of $0.15$ for $Y=0$ and $0.85$ for $Y=1$. If $A$ is forced to be $1$, then $Y$ is 0 when $U_{ay}=U_{by}=0$ and $U_{y}=1$, leading to probabilities of $0.1$ for $Y=0$ and $0.9$ for $Y=1$.
@@
Additionally, both models satisfy monotonicity with respect to $A$ and $B$. There are no cases where $Y_{a=1}=0$ and $Y_{a=0}=1$ in model $M_{1}$, and in model $M_{0}$, the only such case occurs when $Y=B$, which remains unchanged if $A$ is set to 0. Similarly, there are no cases where $Y_{b=1}=0$ and $Y_{b=0}=1$ due to the nature of $Y$ in both models.
@@
When $A$ is set to 0 regardless of $B$ in both models $M_{0}$ and $M_{1}$, only one scenario results in $Y=0$: when $U_{ay}=U_{by}=0$ and $U_{y}\neq 0$. This yields $\Pr_{M_{0}}(Y_{A=0}=0)=\Pr_{M_{1}}(Y_{A=0}=0)=0.15$. However, when $A$ is forced to be $1$, in model $M_{0}$, $Y$ consistently equals $1$ across all scenarios. Yet, in model $M_{1}$, there’s a case where $Y=0$ when $U_{ay}=U_{by}=0$ and $U_{y}=1$, with a probability of $0.1$. Thus, $\Pr_{M_{0}}(Y_{A=1}=1)=1$, different from $\Pr_{M_{1}}(Y_{A=1}=1)=0.9$. Consequently, $ATE_{M_{0}}=0.15$, in contrast to $ATE_{M_{1}}=0.05$. This disparity demonstrates that both $\Pr(Y_{A=1})$ and the ATE lack point identifiability.
Table 4: Calculated probabilities for models $M_{0}$ and $M_{1}$ as presented in the counterexample detailed in Theorem B.1. Despite both models being monotonic, they agree on all observational and factorial (2-treatment) quantities. However, they diverge regarding $\Pr(Y_{A=1})$, making not only this probability but also the ATE non-identifiable
@@
[…]
@@
[…]
@@ Appendix C Parametric Identification of Single-Treatment Estimands Using Factorial Data
Theorem C.1 (AS is sufficient for identifiability of the ATE).
@@
Suppose $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta_{2}b_{i}+\beta_{3}a_{i}b_{i}+u_{i}$ with unobserved $u_{i}$ is a model of the real-world, discrete $b_{i}$ and $a_{i}$, and assume statistical dependence between $a_{i}$ and $b_{i}$ and $u_{i}$. Then the expected outcome of an intervention $A=a$ is identifiable using the observational quantity $\Pr(B=b)$ and the factorial distribution $\operatorname{\mathbb{E}}(Y_{a,b})$ by:
@@
Suppose $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta_{2}b_{i}+u_{i}$ with unobserved $u_{i}$ is a model of the real-world and assume statistical dependence between $a_{i}$ and $b_{i}$ and $u_{i}$. Then the ATE of $A$ on $Y$ is identifiable using factorial data by the quantity in equation 4 as that is a case of AS (Theorem C.1). Additionally, linearity is also a case of interactive linearity (Theorem C.2), so single-treatment quantities are identifiable by Equation 5. Finally, observational data alone is insufficient for identifiability given the standard omitted variable bias (unobserved confounding).